OpenStudy (anonymous):

Need to check my answer.. please let me know if this is the correct A ramp that is 10.0m long is used to push crates over the edge of a ship. The ramp is propped up so that it is 5.00m above the deck. If a crate of mass 100kg is pushed all the way up the 10m ramp with a force of 1100N, starting from the bottom of the ramp, what will be the velocity and direction of the crate as it is hurled off the ramp? Ignore friction First i determined the direction of motion θ = arcsin ( 5.00m/10.0m) = 30 degrees

OpenStudy (anonymous):

I figured that for this problem there is no need to break it down into x and y components since all forces and motion are gonna be in the same direction.. Is that right? So F=ma 1100N = (100kg)a a = 11m/s^2 So the velocity will be V^2 = 2(11m/s^2)(10m) V= 14.83m/s at angle 30degrees Anything wrong with this please let me know! Thanks!

OpenStudy (warriorz13):

I have found nothing wrong with your work but I may also be wrong so is it ok if I mention a couple people to check your work to?